# A Dimca-Greuel type inequality for foliations

### v1

@@ Abstract.
Let $\mathcal{F}$ be a holomorphic foliation at $p\in\mathbb{C}^{2}$, and $B$ be a separatrix of $\mathcal{F}$. We prove the following Dimca-Greuel type inequality $\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<4/3$, where $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$ and $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}[[x,y]]$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$. As a consequence, we provide a new proof of the $\frac{4}{3}$-Dimca-Greuel’s conjecture for singularities of irreducible plane curve germs, with foliations ingredients, that differs from those given by Alberich-Carramiñana, Almirón, Blanco, Melle-Hernández and Genzmer-Hernandes but it is in line with the idea developed by Wang.
@@ 1. Introduction
In [6, Question 4.2], Dimca and Greuel present a family of curves $\{C_{m}\}_{m\geq 1}$ such that the quotient $\frac{\mu(C_{m})}{\tau(C_{m})}$ is strictly increasing with limit $4/3$ and they ask if this property is true for any plane curve singularity. The question was affirmatively answered in [1], [10], and [15] for the irreducible case and in [2] for the reduced case.
It is natural to wonder whether this bound also holds if we consider germs of singular foliations instead of plane curve germs. More precisely, if $\mathcal{F}$ is a germ of singular foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$, then the Milnor number of $\mathcal{F}$ is by definition the dimension of $\mathbb{C}[[x,y]]/(P,Q)$ and we denote it by $\mu_{p}(\mathcal{F})$. Moreover if $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices of $\mathcal{F}$ (see [9], [11]), the Tjurina number of $\mathcal{F}$ with respect to $\mathcal{B}_{0}:f(x,y)=0$ is the dimension $\tau_{p}(\mathcal{F},\mathcal{B}_{0}):=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,Q,f)$. If $\mathcal{F}$ is the hamiltonian foliation given by the $1$-form $df$ for some $f\in\mathbb{C}\{x,y\}$, then the ratio
@@
If the foliation $\mathcal{F}$ is of the second type and $\mathcal{B}$ is a balanced divisor of separatrices for $\mathcal{F}$ then, in [8, Proposition 4.1], we have proved the inequality $\mu_{p}(\mathcal{F})\leq 2\tau_{p}(\mathcal{F},\mathcal{B}_{0})$  generalizing to foliations of second type this well-known inequality in the context of germs of reduced plane curves (see for instance [14] where the general case of hypersurfaces is studied). It is natural to wonder if we can improve the bound $2$ for the ratio (1).
In general, the Dimca-Greuel bound does not hold for foliations, as the following example illustrates: if we borrow from [7, Example 6.5] the family of dicritical foliations $\mathcal{F}_{k}$, $k\geq 3$, given by the $1$-form
@@
However, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a separatrix $B$ of $\mathcal{F}$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition), and we put in the denominator $\tau_{p}(\mathcal{F},B)$ then we get a bound of the Dimca-Greuel type for foliations.
@@
Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. Then
$\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<\frac{4}{3}.$
We will prove Theorem A in Section 3.
@@ 2. Multiplicity of a foliation along a divisor of separatrices
Throughout this text $\mathcal{F}$ denotes a germ of a singular (holomorphic or formal) foliation at $(\mathbb{C}^{2},p)$, given in local coordinates $(x,y)$ centered at $p$ by a (holomorphic or formal) 1-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$ are coprime; or by its dual vector field
$v:=-Q(x,y)\frac{\partial}{\partial x}+P(x,y)\frac{\partial}{\partial y}.$
The algebraic multiplicity $\nu_{p}(\mathcal{F})$ of $\mathcal{F}$ is the minimum of the orders $\text{ord}_{p}(P)$ and $\text{ord}_{p}(Q)$.
Remember that a plane curve germ $f(x,y)=0$ is an $\mathcal{F}-$invariant curve if $\omega\wedge df=(f.h)dx\wedge dy$, for some $h(x,y)\in\mathbb{C}[[x,y]]$ and a separatrix of $\mathcal{F}$ is an irreducible $\mathcal{F}$-invariant curve.
Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$ induced by the vector field $v$ and $B$ be a separatrix of $\mathcal{F}$. Let $\gamma:(\mathbb{C},0)\to(\mathbb{C}^{2},p)$ be a primitive parametrization of $B$, we can consider the multiplicity of $\mathcal{F}$ along $B$ at $p$ defined by
@@
where $\theta(t)$ is the unique vector field at $(\mathbb{C},0)$ such that $\gamma_{*}\theta(t)=v\circ\gamma(t)$, see for instance [4, page 159]. If $\omega=P(x,y)dx+Q(x,y)dy$ is a 1-form inducing $\mathcal{F}$ and $\gamma(t)=(x(t),y(t))$, then
@@
By convention we put $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$. In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [13, Equation (2.2), p. 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$). Hence, if we write $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ where $\mathcal{B}_{0}$ and $\mathcal{B}_{\infty}$ are effective divisors we get
@@
Since $\mu(\mathcal{F},\mathcal{B}_{\infty})\geq 1$ we get
@@
where $\gamma(t)=(x(t),y(t))$ is a Puiseux parametrization of $B$. By [7, Proposition 5.7], we get $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(\mathcal{G}_{f},B)$. But, since $B$ is irreducible, after [7, Proposition 4.7], $\mu_{p}(\mathcal{G}_{f},B)=\mu_{p}(B)$ and so that $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(B).$

### v2

@@ Abstract.
Let $\mathcal{F}$ be a holomorphic foliation at $p\in\mathbb{C}^{2}$, and $B$ be a separatrix of $\mathcal{F}$. Under certain conditions on the reduction of singularities of $\mathcal{F}$ and $B$, we prove the Dimca-Greuel type inequality $\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<4/3$, where $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$ and $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}\{x,y\}$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$. As a consequence, we provide a new proof of the $\frac{4}{3}$-Dimca-Greuel conjecture for singularities of irreducible plane curve germs, with foliations ingredients, that differs from those given by Alberich-Carramiñana, Almirón, Blanco, Melle-Hernández and Genzmer-Hernandes but it is in line with the idea developed by Wang.
@@ 1. Introduction
In [7, Question 4.2], Dimca and Greuel present a family of curves $\{C_{m}\}_{m\geq 1}$ such that the quotient $\frac{\mu(C_{m})}{\tau(C_{m})}$ is strictly increasing with limit $4/3$ and they ask if this property is true for any plane curve singularity. The question was affirmatively answered in [1], [11], and [19] for the irreducible case and in [2] for the reduced case.
It is natural to wonder whether this bound also holds if we consider germs of singular foliations instead of plane curve germs. More precisely, if $\mathcal{F}$ is a germ of singular foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$, then the Milnor number of $\mathcal{F}$ is by definition the dimension of $\mathbb{C}[[x,y]]/(P,Q)$ and we denote it by $\mu_{p}(\mathcal{F})$. The algebraic multiplicity $\nu_{p}(\mathcal{F})$ is the minimum of the orders $\text{ord}_{p}(P)$ and $\text{ord}_{p}(Q)$. Moreover, if $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices of $\mathcal{F}$ (see [10], [12]), the Tjurina number of $\mathcal{F}$ with respect to $\mathcal{B}_{0}:f(x,y)=0$ is the dimension $\tau_{p}(\mathcal{F},\mathcal{B}_{0}):=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,Q,f)$. If $\mathcal{F}$ is the hamiltonian foliation given by the $1$-form $df$ for some $f\in\mathbb{C}\{x,y\}$, then the ratio
@@
If the foliation $\mathcal{F}$ is of the second type (see [10, Definition 2.5]) and $\mathcal{B}$ is a balanced divisor of separatrices for $\mathcal{F}$ then, in [9, Proposition 4.1], we have proved the inequality $\mu_{p}(\mathcal{F})\leq 2\tau_{p}(\mathcal{F},\mathcal{B}_{0})$  generalizing to foliations of second type this well-known inequality in the context of germs of reduced plane curves (see for instance [15] where the general case of hypersurfaces is studied). It is natural to wonder if we can improve the bound $2$ for the ratio (1).
In general, the Dimca-Greuel bound does not hold for foliations, as the following example illustrates: if we borrow from [8, Example 6.5] the family of dicritical foliations $\mathcal{F}_{k}$, $k\geq 3$, $\0\lambda\in\mathbb{C}\backslash\{0\}$, given by the $1$-form
@@
However, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a separatrix $B$ of $\mathcal{F}$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition), and we put in the denominator $\tau_{p}(\mathcal{F},B)$, then, under certain conditions on the reduction of singularities of $\mathcal{F}$ and $B$, we will obtain a bound of the Dimca-Greuel type for foliations. More specifically, in Foliation theory, two type of separatrices of $\mathcal{F}$ are distinguished: isolated separatrices $\text{ISO}_{p}(\mathcal{F})$ and dicritical separatrices $\text{Dic}_{p}(\mathcal{F})$. Isolated separatrices of $\mathcal{F}$ correspond to separatrices that are images, by the reduction process of singularities $\pi$ of $\mathcal{F}$, of separatrices of non-degenerate singularities, or weak or strong separatrices of saddle-nodes singularities in a reduction of singularities of $\mathcal{F}$. On the other hand, dicritical separatrices correspond to separatrices that are images, by $\pi$, of separatrices that intersect transversally with a dicritical divisor in the last step of the reduction of singularities of $\mathcal{F}$. For more details about these definitions, see Section 3.
In order to establish an inequality of Dimca-Greuel type, it is necessary to introduce a number $\text{DG}_{p}(\mathcal{F},B)$ that depends on the reduction of singularities of the pair $(\mathcal{F},B)$ and on the type of separatrix that $B$ is, see Definition 21. We will see that the negativity of this number is crucial in order to obtain an inequality of the Dimca-Greuel type for foliations.
@@
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. If $\text{DG}_{p}(\mathcal{F},B)<0$, then
(2) $\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<\frac{4}{3}.$
We will prove Theorem A in Section 4. In Example 4.3, we will show that we cannot removed the hypothesis over $\text{DG}_{p}(\mathcal{F},B)$.
@@ 2. Multiplicity of a foliation along a divisor of separatrices
Throughout this note $\mathcal{F}$ denotes a germ of a singular (holomorphic or formal) foliation at $(\mathbb{C}^{2},p)$, given in local coordinates $(x,y)$ centered at $p$ by a (holomorphic or formal) 1-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$ are coprime; or by its dual vector field
(3) $v:=-Q(x,y)\frac{\partial}{\partial x}+P(x,y)\frac{\partial}{\partial y}.$
Remember that a plane curve germ $f(x,y)=0$ is an $\mathcal{F}-$invariant curve if $\omega\wedge df=(f.h)dx\wedge dy$, for some $h(x,y)\in\mathbb{C}[[x,y]]$ and a separatrix of $\mathcal{F}$ is an irreducible $\mathcal{F}$-invariant curve. We denote by $\text{Sep}_{p}(\mathcal{F})$ the set of all separatrices of $\mathcal{F}$ at $p$.
Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$ induced by a vector field $v$ as in (3) and $B\in\text{Sep}_{p}(\mathcal{F})$. Let $\gamma:(\mathbb{C},0)\to(\mathbb{C}^{2},p)$ be a primitive parametrization of $B$, we can consider the multiplicity of $\mathcal{F}$ along $B$ at $p$ defined by
@@
where $\theta(t)$ is the unique vector field at $(\mathbb{C},0)$ such that $\gamma_{*}\theta(t)=v\circ\gamma(t)$, see for instance [5, p. 159]. If $\omega=P(x,y)dx+Q(x,y)dy$ is a 1-form inducing $\mathcal{F}$ and $\gamma(t)=(x(t),y(t))$, then
@@
By convention we put $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$. In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [14, Equation (2.2), p. 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$). Hence, if we write $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ where $\mathcal{B}_{0}$ and $\mathcal{B}_{\infty}$ are effective divisors we get
@@
Since $\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})\geq 1$ we have
@@
where $\gamma(t)=(x(t),y(t))$ is a Puiseux parametrization of $B$. By [8, Proposition 5.7], we get $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(\mathcal{G}_{f},B)$. But, since $B$ is irreducible, after [8, Proposition 4.7], $\mu_{p}(\mathcal{G}_{f},B)=\mu_{p}(B)$ and so that $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(B).$ Hence, according to [8, Proposition 6.2], we obtain
(10) $\mu_{p}(\mathcal{F},B)-\tau_{p}(\mathcal{F},B)=\mu_{p}(B)-\tau_{p}(B).$
Consequently when $B$ is quasihomogeneous by ([17], Satz p.123 a) and d)] and (10) or by [20, Theorem 4], $B$ is analytically equivalent to $y^{n}+x^{m}=0$ for some natural coprime numbers $n$, $m>1$, in this case, we get $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$.
3. Reduction of singularities of foliations and their separatrices
We say that $p\in\mathbb{C}^{2}$ is a reduced singularity for $\mathcal{F}$ if the linear part $\text{D}v(p)$ of the vector field $v$ in (3) is non-zero and has eigenvalues $\lambda_{1},\lambda_{2}\in\mathbb{C}$ fitting in one of the cases:
  1. (i)
    $\lambda_{1}\lambda_{2}\neq 0$ and $\lambda_{1}\lambda_{2}\not\in\mathbb{Q}^{+}$ (non-degenerate);
  2. (ii)
    $\lambda_{1}\neq 0$ and $\lambda_{2}\neq 0$ (saddle-node singularity).
In the case $(i)$, there is a system of coordinates $(x,y)$ in which $\mathcal{F}$ is defined by the equation
(11) $\omega=x(\lambda_{1}+a(x,y))dy-y(\lambda_{2}+b(x,y))dx,$
where $a(x,y),b(x,y)\in{\mathbb{C}}[[x,y]]$ are non-units, so that $\text{Sep}_{p}(\mathcal{F})$ is formed by two transversal analytic branches given by $\{x=0\}$ and $\{y=0\}$. In the case $(ii)$, up to a formal change of coordinates, the saddle-node singularity is given by a 1-form of the type
(12) $\omega=x^{k+1}dy-y(1+\lambda x^{k})dx,$
where $\lambda\in\mathbb{C}$ and $k\in\mathbb{Z}^{+}$ are invariants after formal changes of coordinates (see [16, Proposition 4.3]). The curve $\{x=0\}$ is an analytic separatrix, called strong separatrix, whereas $\{y=0\}$ corresponds to a possibly formal separatrix, called weak separatrix. The integer $k+1>1$ is called tangency index of $\mathcal{F}$ with respect to the weak separatrix.
Let $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ be a composition of blow-ups maps. The divisor $\mathcal{D}$ is a finite union of components which are embedding projective lines, crossing normally at corners. If $\mathcal{F}$ is the foliation given by the 1-form $\omega$, we denote ~$\tilde{\mathcal{F}}=\pi^{*}\mathcal{F}$ the strict transform of $\mathcal{F}$, the germ of foliation on $(\tilde{X},\mathcal{D})$ defined locally by $\pi^{*}(\omega)$, after cancelling one-dimensional singular components. For a uniform analysis, we include the possibility of $\pi$ being the identity map and, abusing notation, we set in this case $\tilde{X}=\mathbb{C}^{2}$, $\mathcal{D}=\{p\}$ and $\tilde{\mathcal{F}}=\mathcal{F}$.
With respect to the divisor $\mathcal{D}$, the foliation $\tilde{\mathcal{F}}$ at a point $q\in\mathcal{D}$ can be:
  • Regular, if there are local analytic coordinates $(x,y)$ at $q$ such that $\mathcal{D}\subset\{xy=0\}$ and ~$\tilde{\mathcal{F}}:dx=0$.
  • Singular; if it is not regular.
  • Reduced, if $q$ is a reduced singularity for $\tilde{\mathcal{F}}$ and $\mathcal{D}\subset\text{Sep}_{q}(\tilde{\mathcal{F}})$.
When $\mathcal{D}=\{p\}$, these notions coincide with the ordinary concepts of regular point, singular point and reduced singularity. We say that $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ is a reduction of singularities for $\mathcal{F}$ if all points $q\in\mathcal{D}$ are either regular or reduced singularities. There always exists a reduction of singularities [18]. Besides, there exists a minimal one, in the sense that it factorizes any other reduction of singularities by an additional sequence of blow-ups. All along this note, reductions of singularities are supposed to be minimal.
Given a germ of foliation $\mathcal{F}$ at $(\mathbb{C}^{2},p)$ with a fixed reduction process $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},0)$ for $\mathcal{F}$, a component $D\subset\mathcal{D}$ can be:
  • Non-dicritical, if $D$ is $\tilde{\mathcal{F}}$-invariant. In this case $D$ contains a finite number of reduced singularities. Each non-corner singularity carries a separatrix transversal to $D$, whose projection by $\pi$ is a curve in $\text{Sep}_{p}(\mathcal{F})$.
  • Dicritical, if $D$ is not $\tilde{\mathcal{F}}$-invariant. The definition of reduction of singularities gives that $D$ may intersect only non-dicritical components and that $\tilde{\mathcal{F}}$ is everywhere transverse to $D$. The $\pi$-image of a local leaf of $\mathcal{F}$ at each non-corner point of $D$ belongs to $\text{Sep}_{p}(\mathcal{F})$.
Denote by $\text{Sep}_{p}(D)\subset\text{Sep}_{p}(\mathcal{F})$ the set of separatrices whose strict transforms by $\pi$ intersect the component $D\subset\mathcal{D}$. If $B\in\text{Sep}_{p}(D)$ with $D$ non-dicritical, $B$ is said to be isolated. Otherwise, it is said to be a dicritical separatrix. This engenders the decomposition $\text{Sep}_{p}(\mathcal{F})=\text{ISO}_{p}(\mathcal{F})\cup\text{Dic}_{p}(\mathcal{F})$, where notation is self evident. The set $\text{ISO}_{p}(\mathcal{F})$ is finite and contains all purely formal separatrices. It subdivides further in two classes: weak separatrices – those arising from the weak separatrices of saddle-nodes – and strong separatrices – corresponding to strong separatrices of saddle-nodes and separatrices of non-degenerate singularities. On the other hand, if $\text{Dic}_{p}(\mathcal{F})$ is a non-empty set then it is an infinite set of analytic separatrices of $\mathcal{F}$. A foliation $\mathcal{F}$ is said to be dicritical when $\text{Sep}_{p}(\mathcal{F})$ is infinite, which is equivalent to saying that $\text{Dic}_{p}(\mathcal{F})$ is non-empty. Otherwise, $\mathcal{F}$ is called non-dicritical. When we deal with a simple blow-up $\sigma$ at $p\in\mathbb{C}^{2}$, we say that $\sigma$ is dicritical with respect to $\mathcal{F}$, if the exceptional divisor $\sigma^{-1}(p)$ is not $\tilde{\mathcal{F}}$-invariant. Otherwise, $\sigma$ is called non-dicritical.
Let $\mathcal{F}$ be a holomorphic foliation at $(\mathbb{C}^{2},p)$ and $B\in\text{Sep}_{p}(\mathcal{F})$ On a minimal reduction of singularities $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ of $\mathcal{F}$ we mark the reduction process of singularities of the pair $(\mathcal{F},B)$:
(13) $(\mathcal{F}^{(N)},B^{(N)})\overset{\sigma_{N}}{\longrightarrow}\cdots\overset{\sigma_{3}}{\longrightarrow}(\mathcal{F}^{(2)},B^{(2)})\overset{\sigma_{2}}{\longrightarrow}(\mathcal{F}^{(1)},B^{(1)})\overset{\sigma_{1}}{\longrightarrow}(\mathcal{F}^{(0)},B^{(0)})=(\mathcal{F},B),$
where $\pi=\sigma_{1}\circ\ldots\circ\sigma_{N}$, $B^{(j+1)}$ denotes the strict transform of $B^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $B^{(N)}$ is the normalization of $B$ (or a strict transform of it), $\mathcal{F}^{(j+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in B^{(N)}\cap\mathcal{D}$.
Let
(14) $\alpha_{j}:=\left\{\begin{array}[]{ll}\dim\frac{(\sigma_{j+2}\circ\ldots\circ\sigma_{N})^{*}\Omega_{B^{(j+1)}}}{(\sigma_{j+1}\circ\ldots\circ\sigma_{N})^{*}\Omega_{B^{(j)}}},&\hbox{ if $j=0,\ldots,N-2$}\\ \dim\frac{\Omega_{B^{(N)}}}{\sigma_{N}^{*}\Omega_{B^{(N-1)}}},&\hbox{ if $j=N-1$}\end{array}\right.$
where for a curve $B$, $\Omega_{B}$ denotes the module of Kähler differentials on $B$. We note that $\alpha_{0}$ appears as a term in [19, Equation 4]. Additionally, notice that $\alpha_{j}$ is the $\alpha_{0}$ for $B^{(j)}$. Moreover, we observe that $\alpha_{j}\geq 0$, and the equality holds when the curve $B^{(j)}$ is smooth.
We set
(15) $R_{p}(\mathcal{F},B):=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
where
(16) $m_{q^{(j)}}(\mathcal{F}^{(j)})=\left\{\begin{array}[]{ll}\nu_{q^{(j)}}(\mathcal{F}^{(j)})+1&\hbox{\rm if $\sigma_{j+1}$ is dicritical}\\ \nu_{q^{(j)}}(\mathcal{F}^{(j)})&\hbox{\rm if $\sigma_{j+1}$ is non-dicritical}.\\ \end{array}\right.$
Notice that for $j=0$ we get $q^{(0)}=p$, $\mathcal{F}^{(0)}=\mathcal{F}$ and the equality (16) for $j=0$ becomes:
(17) $m_{p}(\mathcal{F})=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{F})+1&\hbox{\rm if$\sigma$ is dicritical}\\ \nu_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is non-dicritical},\\ \end{array}\right.$
where $\sigma$ is a blow-up at the point $p$.
The following lemma relates $R_{p}(\mathcal{F},B)$ to the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ (respectively $\mathcal{F}^{(N)}$) with respect to $B$ (respectively $B^{(N)}$).
Lemma 3.1.
With the above notations we get
$R_{p}(\mathcal{F},B)=GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})-GSV_{p}(\mathcal{F},B)+\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right].$
Proof.
From blow-up formula of the GSV index [4, p.30] we get
(18) $GSV_{q^{(j+1)}}(\mathcal{F}^{(j+1)},B^{(j+1)})$ $=$ $GSV_{q^{(j)}}(\mathcal{F}^{(j)},B^{(j)})$
$+$ $\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})),\,\,\,\,\hbox{\rm for }\,\,\,0\leq j\leq N-1.$
Up substituting the equality (18) in (15), we get
$R_{p}(\mathcal{F},B)$ $=$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right]$
$=$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)}))\right]$
$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right]$
$=$ $\sum_{j=0}^{N-1}\left[GSV_{q^{(j+1)}}(\mathcal{F}^{(j+1)},B^{(j+1)})-GSV_{q^{(j)}}(\mathcal{F}^{(j)},B^{(j)})\right]$
$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right]$
$=$ $GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})-GSV_{p}(\mathcal{F},B)$
$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right],$
and the lemma follows. ∎
For a germ of a holomorphic foliation $\mathcal{F}$ and a separatrix $B\in\text{Sep}_{p}(\mathcal{F})$ at $(\mathbb{C}^{2},p)$, we introduce the number $\text{DG}_{p}(\mathcal{F},B)$ as follows:
(21) $\text{DG}_{p}(\mathcal{F},B)$ ::=:= $\left\{\begin{array}[]{ll}R_{p}(\mathcal{F},B)-\tau_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})&\hbox{\rm if $B\in\text{ISO}_{p}(\mathcal{F})$}\\ R_{p}(\mathcal{F},B)&\hbox{\rm if $B\in\text{Dic}_{p}(\mathcal{F})$}.\\ \end{array}\right.$
From (14), we deduce that $\alpha_{i}$ are analytic invariants, thus we have $\text{DG}_{p}(\mathcal{F},B)$ is also an analytic invariant depending of $\mathcal{F}$ and $B$. As a consequence of Lemma 3.1, we can establish $\text{DG}_{p}(\mathcal{F},B)$ in terms of the GSV-indices of the foliations $\mathcal{F}$ and $\mathcal{F}^{(N)}$ and some numbers depending of the resolution of the singularities of $B$.